# Projection-free computation of robust controllable sets with constrained zonotopes

### v1

@@ Abstract
We study the problem of computing robust controllable sets for discrete-time linear systems with additive uncertainty. We propose a tractable and scalable approach to inner- and outer-approximate robust controllable sets using constrained zonotopes, when the additive uncertainty set is a symmetric, convex, and compact set. Our least-squares-based approach uses novel closed-form approximations of the Pontryagin difference between a constrained zonotopic minuend and a symmetric, convex, and compact subtrahend. Unlike existing approaches, our approach does not rely on convex optimization solvers, and is projection-free for ellipsoidal and zonotopic uncertainty sets. We also propose a least-squares-based approach to compute a convex, polyhedral outer-approximation to constrained zonotopes, and characterize sufficient conditions under which all these approximations are exact. We demonstrate the computational efficiency and scalability of our approach in several case studies, including the design of abort-safe rendezvous trajectories for a spacecraft in near-rectilinear halo orbit under uncertainty. Our approach can inner-approximate a 20-step robust controllable set for a 100-dimensional linear system in under 15 seconds on a standard computer.
@@ 1 Introduction
Robust controllable (RC) sets characterize a set of states from which a collection of possibly time-varying state constraints can be satisfied by a controlled state trajectory, under bounded control authority and uncertainty. RC sets are essential for robust model predictive control [1, 2, 3, 4, 5], fault-tolerant control [6, 7, 8], and verification of dynamical systems [9, 10, 11, 12], and have been used in a broad range of applications in space [13, 11, 10, 12], transportation [14], and robotics [15, 16]. For discrete-time linear systems with additive uncertainty and bounded, polytopic state and input constraints, exact RC sets are known to be polytopes, and the RC sets may be computed using set computations on polytopes. However, polytope-based RC set computation involves projection that has a combinatorial computational complexity and causes numerical issues for high-dimensional systems and/or over long horizons [15, 3, 17]. In this paper, we focus in addressing these shortcomings, and propose novel theory and algorithms for efficient and scalable computation of inner- and outer-approximations of RC sets, that admit closed-form description of the sets involved and can be computed without relying on convex optimization solvers in several cases.
We are motivated by the problem of designing active, abort-safe, spacecraft rendezvous trajectories with the Lunar gateway [18]. In abort-safe rendezvous, we seek rendezvous trajectories that nominally steer the control-constrained spacecraft towards the rendezvous target, while allowing for the possibility of diverting away from the rendezvous target in the event of failure. Often, failures in space applications are characterized by partial or complete loss of actuation, and increased actuation and navigational uncertainties that may appear as additive uncertainties in the dynamics. For designing such rendezvous trajectories, we use RC sets to characterize a set of state constraints that a nominal trajectory should satisfy, where the RC sets encode the desirable property of safe abort under limited available actuation and uncertainty, similarly to [13, 11]. However, polytope-based computation of RC sets for high-dimensional systems suffers from numerical issues, primarily due to projection [15, 3, 17]. On the other hand, our approach can efficiently compute the required high-dimensional RC sets.
@@
The main contributions of this paper are as follows: 1) a tractable inner- and outer-approximation of the Pontryagin difference between a constrained zonotopic minuend and a symmetric, convex, and compact subtrahend, 2) an inner- and outer-approximation of the RC sets using the proposed approximations to the Pontryagin difference, 3) a closed-form description of a polyhedral outer-approximation of a constrained zonotope, and 4) sufficient conditions under which the approximations proposed above are exact. All of our approaches utilize least-squares method as the primary tool for which efficient implementations are widely available [20], and does not typically rely on convex optimization solvers as required in [9]. Additionally, the proposed approximations to the Pontryagin difference admit closed-form expressions when the subtrahend is an ellipsoid, a zonotope, or a convex hull of a collection of symmetric intervals. These features together enable an inner-approximation of a 20-step RC set for a 100-dimensional linear system in less than 15 seconds on a standard computer.
The rest of this paper is organized as follows: Section 2 provides the necessary mathematical background and states the problem statements of interest. Sections 3 and 4 describe the proposed approaches to approximate the Pontryagin difference between a constrained zonotopic minuend and a symmetric, convex, and compact subtrahend respectively. Section 5 applies these approaches to the computation of RC sets, and Section 6 presents several case studies that demonstrate the utility and scalability of the proposed approach. Section 7 applies the method to the (simplified) motivating problem of active abort-safe rendezvous with the Lunar gateway. Section 8 summarizes the paper.
@@ 2 Preliminaries
$0_{n\times m}$ and $1_{n\times m}$ are matrices of zeros and ones in $\mathbb{R}^{n\times m}$ respectively, $I_{n}$ is the $n$-dimensional identity matrix, $\mathbb{N}_{[a:b]}$ is the subset of natural numbers between (and including) $a,b\in\mathbb{N}$, $a\leq b$, $e_{i}$ is the standard axis vectors of $\mathbb{R}^{n}$, and ${\|\cdot\|}_{p}$ is the $\ell_{p}$-norm of a vector. Let $M$ be a matrix and $M_{1}$ ($M_{2}$, resp.) be another matrix of the same height (width, resp.) as $M$. Then, $[M,M_{1}]$ ($[M;M_{2}]$, resp.) denotes the matrix obtained by concatenating $M$ and $M_{1}$ horizontally (concatenating $M$ and $M_{2}$ vertically, resp.). For a matrix $M\in\mathbb{R}^{m\times n}$ with full row rank, $M^{\dagger}=M^{\top}{\left({MM^{\top}}\right)}^{-1}$ denotes its (right) pseudoinverse, and $x=M^{\dagger}v$ solves the system of linear equation $Mx=v$ for any vector $v\in\mathbb{R}^{m}$ [20, Sec. 11.5]. Given a vector $d\in\mathbb{R}^{n}$, $\operatorname{diag}(d)$ is a $n$-dimensional diagonal matrix with $d_{i}$ as its diagonal entries.
A set $\mathcal{S}\subset\mathbb{R}^{n}$ is said to be symmetric about any $c\in\mathbb{R}^{n}$, if $c+x\in\mathcal{S}$ implies $c-x\in\mathcal{S}$ for any $x\in\mathbb{R}^{n}$. For any set $\mathcal{S}\subseteq\mathbb{R}^{n}$, we denote its convex hull and affine hull by $\operatorname{CH}\left({\mathcal{S}}\right)$ and $\operatorname{AH}\left({\mathcal{S}}\right)$ respectively. Recall that $\operatorname{AH}\left({\mathcal{S}}\right)$ is an affine set such that $\mathcal{S}\subseteq\mathcal{A}\Rightarrow\operatorname{AH}\left({\mathcal{S}}\right)\subseteq\mathcal{A}$ for any affine set $\mathcal{A}$. The affine dimension of a set $\mathcal{S}$ is the dimension of the subspace associated with $\operatorname{AH}\left({\mathcal{S}}\right)$. A full-dimensional set in $\mathbb{R}^{n}$ is a non-empty set with an affine dimension of $n$. See [21, Sec. 2.1] for more details.
@@ 2.1 Set representations
Let $\mathcal{P},\mathcal{C},\mathcal{I},\mathcal{Z}$, and $\mathcal{E}$ be convex and compact sets in $\mathbb{R}^{n}$. Then, $\mathcal{P}$ is a convex polytope if (1a) holds; $\mathcal{C}$ is a constrained zonotope if (1b) holds; $\mathcal{I}$ is a convex union of symmetric intervals if (1c) holds; $\mathcal{Z}$ is a zonotope if (1d) holds; and $\mathcal{E}$ is an ellipsoid if (1e) holds:
$\exists(H_{P},k_{P})\in\mathbb{R}^{N_{P}\times n}\times\mathbb{R}^{N_{P}}$ $:\mathcal{P}=\{x:H_{P}x\leq k_{P}\},$ (1a)
$\exists(G_{C},c_{C},A_{C},b_{C})\in\mathbb{R}_{C}$ $:\mathcal{C}=\left\{G_{C}\xi+c_{C}\middle|{\|\xi\|}_{\infty}\leq 1,\ A_{C}\xi=b_{C}\right\},$ (1b)
$\exists(G_{I},c_{I})\in\mathbb{R}^{n\times N_{I}}\times\mathbb{R}^{n}$ :=$:\mathcal{I}=\{G_{I}\xi+c_{I}:{\|\xi\|}_{1}\leq 1\},$ (1c)
$\exists(G_{Z},c_{Z})\in\mathbb{R}^{n\times N_{Z}}\times\mathbb{R}^{n}$ $:\mathcal{Z}=\{G_{Z}\xi+c_{Z}:{\|\xi\|}_{\infty}\leq 1\},$ (1d)
$\exists(G_{E},c_{E})\in\mathbb{R}^{n\times n}\times\mathbb{R}^{n}$ :=$:\mathcal{E}=\{G_{E}\xi+c_{E}:{\|\xi\|}_{2}\leq 1\},$ (1e)
where $\mathbb{R}_{C}\triangleq\mathbb{R}^{n\times N_{C}}\times\mathbb{R}^{n}\times\mathbb{R}^{M_{C}\times N_{C}}\times\mathbb{R}^{M_{C}}$. We refer to unbounded convex sets in (1a) as convex polyhedra.
Zonotopes $\mathcal{Z}$, ellipsoids $\mathcal{E}$, and convex unions of symmetric intervals $\mathcal{I}$ are affine transformations of unit balls defined using $\ell_{\infty}$-norms, $\ell_{2}$-norms, and $\ell_{1}$-norms respectively, and are symmetric about $c_{Z}$, $c_{E}$, and $c_{I}$ respectively in (1d), (1e), and (1c). Specifically, by the definition of $\ell_{1}$-norm, $\mathcal{I}$ is the convex hull of the union of symmetric intervals defined by the columns of $G_{\mathcal{I}}=[g_{1},g_{2},\ldots,g_{N_{I}}]$,
$\mathcal{I}=c_{I}+\operatorname{CH}\left({\cup_{j\in\mathbb{N}_{[1:N_{I}]}}\llbracket-g_{j},g_{j}\rrbracket}\right),$ (2)
where $\llbracket-g,g\rrbracket=\{g\theta:-1\leq\theta\leq 1\}$ denotes an interval.
From [6, Thm. 1], every convex polytope $\mathcal{P}$ can be represented as a constrained zonotope $\mathcal{C}$, and vice versa. Additionally, [6, Thm. 1] provides a systematic approach to convert $\mathcal{P}$ in the form of (1a) to a constrained zonotope $\mathcal{C}$ in the form of (1b). However, converting $\mathcal{C}$ to $\mathcal{P}$ exactly is known to be a computationally demanding process, with existing approaches applicable only for low-dimensional constrained zonotopes [6, Prop. 3].
From (1b), a constrained zonotope $\mathcal{C}$ can also be interpreted as the image of a linearly-constrained unit-hypercube $\mathcal{B}_{\infty}(A_{C},b_{C})$ under an affine mapping defined by $(G_{C},c_{C})$, i.e., $\mathcal{C}=c_{C}+G_{C}\mathcal{B}_{\infty}(A_{C},b_{C})$ where
$\mathcal{B}_{\infty}(A_{C},b_{C})\triangleq\{\xi:{\|\xi\|}_{\infty}\leq 1,A_{C}\xi=b_{C}\}.$ (3)
In (3), $\mathcal{B}_{\infty}(A_{C},b_{C})$ is the intersection of a unit-hypercube in $\mathbb{R}^{N_{C}}$ and $M_{C}$ linear equalities. By definition, $\mathcal{C}\neq\emptyset$ if and only if $\mathcal{B}_{\infty}(A_{C},b_{C})\neq\emptyset$.
@@
(Representation complexity) The representation complexity of a constrained zonotope $\mathcal{C}$ is $\mathscr{C}(\mathcal{C})=\left({n,N_{C},M_{C}}\right)$.
[…]
@@
$\vartheta_{\mathcal{S}}(\nu)=\arg\sup\nolimits_{s\in\mathcal{S}}\ \nu^{\top}s,\text{ and }\rho_{\mathcal{S}}(\nu)=\nu^{\top}\vartheta_{\mathcal{S}}(\nu).$ (4)
Using dual norms [21, Sec. A.1.6] and properties of support function [22, Prop. 2], we have closed-form expressions for the support function of zonotopes $\mathcal{Z}$, ellipsoids $\mathcal{E}$, and convex unions of symmetric intervals $\mathcal{I}$,
@@ 2.2 Set operations
$R\mathcal{C}$$\triangleq\{Ru:u\in\mathcal{C}\},$ (6a)
$\mathcal{C}\oplus\mathcal{S}$$\triangleq\{u+v:u\in\mathcal{C},\ v\in\mathcal{S}\},$ (6b)
$\mathcal{C}\cap_{R}\mathcal{W}$$\triangleq\{u\in\mathcal{C}:Ru\in\mathcal{W}\},$ (6c)
$\mathcal{C}\ominus\mathcal{S}$$\triangleq\{u:\forall v\in\mathcal{S},u+v\in\mathcal{C}\}.$ (6d)
@@
Constrained zonotopes admit closed-form expressions for various set operations described in (2.2). From [6, Prop. 1] and [7, Sec. 3.2],
@@
$\mathcal{C}\oplus\mathcal{S}$ $=\left([G_{C},G_{S}],c_{C}+c_{S},[A_{C},0;0,A_{S}],[b_{C};b_{S}]\right),$ (7b)
$\mathcal{C}\cap_{R}\mathcal{W}$ $=\left([G_{C},0],c_{C},[A_{C},0;0,A_{W};RG_{C},-G_{W}],[b_{C};b_{W};c_{W}-Rc_{C}]\right),$ (7c)
$\mathcal{C}\cap\mathcal{H}$ $=\left([G_{C},0],c_{C},[A_{C},0;p^{\top}G_{C},d_{m}/2],[b_{C};(q+p^{\top}c_{C}-\|p^{\top}G_{C}\|_{1})/2]\right),$ (7d)
where =$\mathcal{H}=\{x:p^{\top}x\leq q\}\subset\mathbb{R}^{n}$ is an halfspace, and (7d) also enables an exact computation of the intersection of a constrained zonotope and a convex polyhedron.
To the best of our knowledge, no closed-form expression similar to (2.2) is available for the Pontryagin difference (6d). On the other hand, $\mathcal{P}\ominus\mathcal{S}$ for a convex polyhedral minuend $\mathcal{P}$ (1a) defined by $N_{P}$ linear constraints and a convex and compact subtrahend $\mathcal{S}$ is given by
$\mathcal{P}\ominus\mathcal{S}$ =$=\{x:H_{P}x\leq k_{P}-[\rho_{\mathcal{S}}(h_{1});\ \ldots;\rho_{\mathcal{S}}(h_{N_{P}})]\},$ (8)
with $H_{P}=[h_{1}^{\top};h_{2}^{\top};\ldots;h_{N_{P}}^{\top}]$ [23, Thm. 2.3]. A direct use of (8) to implement $\mathcal{C}\ominus\mathcal{S}$ is prevented by the difficulty in converting $\mathcal{C}$ in the form of (1b) to (1a).
@@
minimizeSumAbs(Γ)s. $\begin{array}[]{cl}{\text{minimize}}&\ \ \operatorname{SumAbs}(\Gamma)\\ \text{s.\ t.}&\ \ [G_{C};A_{C}]\Gamma=[G_{S};0_{M_{C}\times N_{S}}],\\ \forall i\in\mathbb{N}_{[1:N_{C}]},&\ \ \sum_{j=1}^{N_{S}}|\Gamma_{ij}|\leq 1,\\ \end{array}\vspace*{-2.5em}$ (12)
@@
The representation complexity of $\mathcal{M}^{-}$ is identical to $\mathcal{C}$, which is a desirable feature when using (14) in set recursions involving the Pontryagin difference. However, solving (12) repeatedly can become computationally expensive for large $N_{C}$ and/or $N_{S}$. Additionally, it is unclear if such an approach extends to subtrahends beyond zonotopes, since it uses polytopic containment results from [24]. We address these shortcomings in this paper.
@@
where $\mathcal{M}^{-}$ is available in closed-form, and the representation complexity of $\mathcal{M}^{-}$ and $\mathcal{C}$ are identical. Also, provide sufficient conditions for $\mathcal{M}^{-}=\mathcal{M}=\mathcal{M}^{+}$.
Our approach to address Problem 1 follows a similar two-stage approach to that in [9] to obtain (14). However, unlike [9], we provide closed-form expressions for $\Gamma$ and $D$ used in (13) and (14) for a broader class of subtrahends that includes ellipsoids and convex unions of symmetric intervals. We also propose a constrained zonotopic outer-approximation to $\mathcal{M}$, and provide sufficient conditions under which these approximations are exact. The proposed outer-approximation relies on solving the following problem.
@@
We will also briefly explore the connections between our proposed optimization-free two-stage approach and the optimization-based two-stage approach in [9].

### v2

@@ Abstract
We study the problem of computing robust controllable sets for discrete-time linear systems with additive uncertainty. We propose a tractable and scalable approach to inner- and outer-approximate robust controllable sets using constrained zonotopes, when the additive uncertainty set is a symmetric, convex, and compact set. Our least-squares-based approach uses novel closed-form approximations of the Pontryagin difference between a constrained zonotopic minuend and a symmetric, convex, and compact subtrahend. We obtain these approximations using two novel canonical representations for full-dimensional constrained zonotopes. Unlike existing approaches, our approach does not rely on convex optimization solvers, and is projection-free for ellipsoidal and zonotopic uncertainty sets. We also propose a least-squares-based approach to compute a convex, polyhedral outer-approximation to constrained zonotopes, and characterize sufficient conditions under which all these approximations are exact. We demonstrate the computational efficiency and scalability of our approach in several case studies, including the design of abort-safe rendezvous trajectories for a spacecraft in near-rectilinear halo orbit under uncertainty. Our approach can inner-approximate a 20-step robust controllable set for a 100-dimensional linear system in under 15 seconds on a standard computer.
@@ 1 Introduction
Robust controllable (RC) sets characterize a set of states from which a collection of possibly time-varying state constraints can be satisfied by a controlled state trajectory, under bounded control authority and uncertainty. RC sets are essential for robust model predictive control [1, 2, 3, 4], fault-tolerant control [5, 6, 7, 8], and verification of dynamical systems [9, 10, 11, 12], and have been used in a broad range of applications in space [13, 11, 10, 12], transportation [14], and robotics [15, 16]. For discrete-time linear systems with additive uncertainty and bounded, polytopic state and input constraints, exact RC sets are known to be polytopes, and the RC sets may be computed using set computations on polytopes. However, polytope-based RC set computation involves projection that has a combinatorial computational complexity and causes numerical issues for high-dimensional systems and/or over long horizons [16, 3, 17, 18]. In this paper, we focus in addressing these shortcomings, and propose novel theory and algorithms for efficient and scalable computation of inner- and outer-approximations of RC sets, that admit closed-form description of the sets involved and can be computed without relying on convex optimization solvers in several cases.
We are motivated by the problem of designing active, abort-safe, spacecraft rendezvous trajectories with the Lunar gateway [19]. In abort-safe rendezvous, we seek rendezvous trajectories that nominally steer the control-constrained spacecraft towards the rendezvous target, while allowing for the possibility of diverting away from the rendezvous target in the event of failure. Often, failures in space applications are characterized by partial or complete loss of actuation, and increased actuation and navigational uncertainties that may appear as additive uncertainties in the dynamics. For designing such rendezvous trajectories, we use RC sets to characterize a set of state constraints that a nominal trajectory should satisfy, where the RC sets encode the desirable property of safe abort under limited available actuation and uncertainty, similarly to [13, 11]. However, polytope-based computation of RC sets for high-dimensional systems suffers from numerical issues, primarily due to projection [16, 3, 18, 17]. On the other hand, our approach can compute the required high-dimensional RC sets.
@@
The main contributions of this paper are as follows: 1) a tractable inner- and outer-approximation of the Pontryagin difference between a constrained zonotopic minuend and a symmetric, convex, and compact subtrahend, 2) an inner- and outer-approximation of the RC sets using the proposed approximations to the Pontryagin difference, 3) a closed-form description of a polyhedral outer-approximation of a constrained zonotope, and 4) sufficient conditions under which the approximations proposed above are exact. We propose two canonical representations for full-dimensional constrained zonotopes, which facilitate the use of the least-squares method as the primary tool for all of our approaches. Efficient implementations for the least-squares method are well-known [21]. Unlike [9], the proposed approximations to the Pontryagin difference admit closed-form expressions when the subtrahend is an ellipsoid, a zonotope, or a convex hull of a collection of symmetric intervals. These features together enable an inner-approximation of a 20-step RC set for a 100-dimensional linear system in less than 15 seconds on a standard computer.
The rest of this paper is organized as follows: Sec. 2 provides the necessary mathematical background and states the problem statements of interest. Sec. 3 and 4 describe the proposed approaches to approximate the Pontryagin difference between a constrained zonotopic minuend and a symmetric, convex, and compact subtrahend. Sec. 5 applies these approaches to the computation of RC sets, and Sec. 6 presents several case studies that demonstrate the utility and scalability of the proposed approach. Sec. 7 applies the method to the (simplified) motivating problem of active abort-safe rendezvous with the Lunar gateway. Sec. 8 summarizes the paper.
@@ 2 Preliminaries
$0_{n\times m}$ and $1_{n\times m}$ are matrices of zeros and ones in $\mathbb{R}^{n\times m}$ respectively, $I_{n}$ is the $n$-dimensional identity matrix, $\mathbb{N}_{[a:b]}$ is the subset of natural numbers between (and including) $a,b\in\mathbb{N}$, $a\leq b$, $\llbracket a,b\rrbracket=\{x\in\mathbb{R}\mid a\leq x\leq b\}$, $e_{i}$ is the standard axis vectors of $\mathbb{R}^{n}$, and ${\|\cdot\|}_{p}$ is the $\ell_{p}$-norm of a vector. Let $M$ be a matrix and $M_{1}$ ($M_{2}$, resp.) be another matrix of the same height (width, resp.) as $M$. Then, $[M,M_{1}]$ ($[M;M_{2}]$, resp.) denotes the matrix obtained by concatenating $M$ and $M_{1}$ horizontally (concatenating $M$ and $M_{2}$ vertically, resp.). For a matrix $M\in\mathbb{R}^{m\times n}$ with full row rank, $M^{\dagger}=M^{\top}{\left({MM^{\top}}\right)}^{-1}$ denotes its (right) pseudoinverse, and $x=M^{\dagger}v$ solves the system of linear equation $Mx=v$ for any vector $v\in\mathbb{R}^{m}$ [21, Sec. 11.5]. Given $d\in\mathbb{R}^{n}$, $\operatorname{diag}(d)$ is a $n$-dimensional diagonal matrix with diagonal entries $d_{i}$.
A set $\mathcal{S}\subset\mathbb{R}^{n}$ is said to be symmetric about any $c\in\mathbb{R}^{n}$, if $c+x\in\mathcal{S}$ implies $c-x\in\mathcal{S}$ for any $x\in\mathbb{R}^{n}$. For any set $\mathcal{S}\subseteq\mathbb{R}^{n}$, we denote its convex hull and affine hull by $\operatorname{CH}\left({\mathcal{S}}\right)$ and $\operatorname{AH}\left({\mathcal{S}}\right)$ respectively. Recall that $\operatorname{AH}\left({\mathcal{S}}\right)$ is an affine set such that $\mathcal{S}\subseteq\mathcal{A}\Rightarrow\operatorname{AH}\left({\mathcal{S}}\right)\subseteq\mathcal{A}$ for any affine set $\mathcal{A}$. The affine dimension of a set $\mathcal{S}$ is the dimension of the subspace associated with $\operatorname{AH}\left({\mathcal{S}}\right)$. A full-dimensional set in $\mathbb{R}^{n}$ is a non-empty set with an affine dimension of $n$. See [22, Sec. 2.1] for more details.
@@ 2.1 Set representations
Let $\mathcal{C}$ be a convex and compact polytope in $\mathbb{R}^{n}$. We consider two representations of $\mathcal{C}$H-Rep polytope (1a) and constrained zonotope (1b),
$\mathcal{C}$ =$=\{x\mid H_{C}x\leq k_{C}\},$ (1a)
$\mathcal{C}$ =$=\left\{G_{C}\xi+c_{C}\ \middle|\ {\|\xi\|}_{\infty}\leq 1,\ A_{C}\xi=b_{C}\right\},$ (1b)
with $H_{C}\in\mathbb{R}^{{L_{C}}\times n}$, $k_{C}\in\mathbb{R}^{L_{C}}$, $G_{C}\in\mathbb{R}^{n\times{N_{C}}}$, $c_{C}\in\mathbb{R}^{n}$, $A_{C}\in\mathbb{R}^{{M_{C}}\times{N_{C}}}$, and $b_{C}\in\mathbb{R}^{{M_{C}}}$. Here, (1a) is the intersection of ${L_{C}}$ halfspaces and (1b) is an affine transformation of $\mathcal{B}_{\infty}(A_{C},b_{C})$, $\mathcal{C}=c_{C}+G_{C}\mathcal{B}_{\infty}(A_{C},b_{C})$ with
$\mathcal{B}_{\infty}(A_{C},b_{C})\triangleq\{\xi\mid{\|\xi\|}_{\infty}\leq 1,A_{C}\xi=b_{C}\}.$ (2)
In (2), $\mathcal{B}_{\infty}(A_{C},b_{C})$ is the intersection of a unit-hypercube in $\mathbb{R}^{{N_{C}}}$ and ${M_{C}}$ linear equalities. By definition, $\mathcal{C}\neq\emptyset$ if and only if $\mathcal{B}_{\infty}(A_{C},b_{C})\neq\emptyset$. Other respresentations of $\mathcal{C}$, apart from (1), include vertex representation (V-Rep polytope) [3] and AH-polytopes [23]. We refer to unbounded H-Rep polytopes as convex polyhedra, and use $\mathcal{C}=(G_{C},c_{C},A_{C},b_{C})$ to denote a polytope $\mathcal{C}$ in constrained zonotope representation (1b).
The equivalence of the representations in (1) was established in [5, Thm. 1]. Additionally, [5, Thm. 1] provides a systematic approach to convert H-Rep polytopes (1a) to constrained zonotopes (1b). However, an exact conversion of (1b) to (1a) is known to be computationally demanding, with existing approaches applicable only for low-dimensional constrained zonotopes [5, Prop. 3].
The representations (1) are not unique. For H-Rep polytopes, a canonical reduction (up to a permutation of rows) is available using linear programming [3, 24]. However, no such canonical reduction is available for constrained zonotopes, to the best of our knowledge. On the other hand, several exact and approximate techniques are available to reduce the representation complexity of a given constrained zonotope [5, 6, 25].
@@
(Representation complexity) [5] Let $\mathscr{O}_{{C}}\triangleq(N_{C}-M_{C})/n$ be the degrees-of-freedom order of a constrained zonotope $\mathcal{C}=(G_{C},c_{C},A_{C},b_{C})$. Then, the representation complexity of $\mathcal{C}$ is $\mathscr{C}(\mathcal{C})=(M_{C},\mathscr{O}_{{C}})$.
Zonotopes $\mathcal{Z}$, ellipsoids $\mathcal{E}$, and convex unions of symmetric intervals $\mathcal{I}$ are affine transformations of unit balls defined using $\ell_{\infty}$-norms, $\ell_{2}$-norms, and $\ell_{1}$-norms respectively. Formally, we define $\mathcal{Z},\mathcal{E},\mathcal{I}\subset\mathbb{R}^{n}$ as follows,
$\mathcal{Z}$$\triangleq\{G_{Z}\xi+c_{Z}\mid{\|\xi\|}_{\infty}\leq 1\},$ (3a)
$\mathcal{E}$$\triangleq\{G_{E}\xi+c_{E}\mid{\|\xi\|}_{2}\leq 1\},$ (3b)
$\mathcal{I}$$\triangleq\{G_{I}\xi+c_{I}\mid{\|\xi\|}_{1}\leq 1\},$ (3c)
with $G_{Z}\in\mathbb{R}^{n\times{N_{Z}}}$, $G_{E}\in\mathbb{R}^{n\times n}$, $G_{I}\in\mathbb{R}^{n\times{N_{I}}}$, and $c_{Z},c_{E},c_{I}\in\mathbb{R}^{n}$. The sets $\mathcal{Z},\mathcal{E},\mathcal{I}$ are symmetric about $c_{Z},c_{E},c_{I}$. We denote these sets using $(G,c)$.
@@
$\vartheta_{\mathcal{S}}(\nu)\triangleq\arg\sup\nolimits_{s\in\mathcal{S}}\ \nu^{\top}s,\text{ and }\rho_{\mathcal{S}}(\nu)\triangleq\nu^{\top}\vartheta_{\mathcal{S}}(\nu).$ (4)
Using dual norms [22, Sec. A.1.6] and properties of support function [26, Prop. 2], we have closed-form expressions for the support function of zonotopes $\mathcal{Z}$, ellipsoids $\mathcal{E}$, and convex unions of symmetric intervals $\mathcal{I}$,
@@ 2.2 Set operations
$R\mathcal{C}$$\triangleq\{Ru\mid u\in\mathcal{C}\},$ (6a)
$\mathcal{C}\oplus\mathcal{S}$$\triangleq\{u+v\mid u\in\mathcal{C},\ v\in\mathcal{S}\},$ (6b)
$\mathcal{C}\cap_{R}\mathcal{W}$$\triangleq\{u\in\mathcal{C}\mid Ru\in\mathcal{W}\},$ (6c)
$\mathcal{C}\ominus\mathcal{S}$$\triangleq\{u\mid\forall v\in\mathcal{S},u+v\in\mathcal{C}\}.$ (6d)
@@
Constrained zonotopes admit closed-form expressions for various set operations described in (6). From [5, 6],
@@
$\mathcal{C}\oplus\mathcal{S}$ $=\left([G_{C},G_{S}],c_{C}+c_{S},[A_{C},0;0,A_{S}],\right.$
$\qquad\left.[b_{C};b_{S}]\right),$ (7b)
$\mathcal{C}\cap_{R}\mathcal{W}$ $=\left([G_{C},0],c_{C},[A_{C},0;0,A_{W};RG_{C},-G_{W}],\right.$
$\qquad\left.[b_{C};b_{W};c_{W}-Rc_{C}]\right),$ (7c)
$\mathcal{C}\cap\mathcal{H}$ $=\left([G_{C},0],c_{C},[A_{C},0;p^{\top}G_{C},d_{m}/2],\right.$
$\qquad\left.[b_{C};(q+p^{\top}c_{C}-\|p^{\top}G_{C}\|_{1})/2]\right),$ (7d)
where =$\mathcal{H}=\{x\mid p^{\top}x\leq q\}\subset\mathbb{R}^{n}$ is an halfspace, and (7d) also enables an exact computation of the intersection of a constrained zonotope and a convex polyhedron.
To the best of our knowledge, the Pontryagin difference (6d) involving a constrained zonotopic minuend does not have a closed-form expression, similar to (7[6, 9]. In fact, given a constrained zonotope $\mathcal{C}$ and a zonotope $\mathcal{Z}$, it is impossible to find a polynomial-size constrained zonotope $\mathcal{C}\ominus\mathcal{Z}$ in polynomial-time, unless P=NP [9, Prop. 1]. On the other hand, given a H-Rep polytope $\mathcal{P}=\{x\mid H_{P}x\leq k_{P}\}$ and a convex and compact subtrahend $\mathcal{S}$, a H-Rep polytope $\mathcal{P}\ominus\mathcal{S}$ is available in closed-form,
$\mathcal{P}\ominus\mathcal{S}$ =$=\{x\mid H_{P}x\leq k_{P}-[\rho_{\mathcal{S}}(h_{1});\ \ldots;\rho_{\mathcal{S}}(h_{{L_{P}}})]\},$ (8)
with $H_{P}=[h_{1}^{\top};h_{2}^{\top};\ldots;h_{{L_{C}}}^{\top}]$ [27, Thm. 2.3]. A direct use of (8) to compute a constrained zonotope $\mathcal{C}\ominus\mathcal{S}$ for a constrained zonotope minuend $\mathcal{C}$ is prevented by the difficulty in converting (1b) to (1a).
@@
minimizeSumAbs(Γ)s. $\begin{array}[]{cl}{\text{minimize}}&\ \ \operatorname{SumAbs}(\Gamma)\\ \text{s.\ t.}&\ \ [G_{C};A_{C}]\Gamma=[G_{S};0_{{M_{C}}\times{N_{S}}}],\\ \forall i\in\mathbb{N}_{[1:{N_{C}}]},&\ \ \sum_{j=1}^{{N_{S}}}|\Gamma_{ij}|\leq 1,\\ \end{array}$ (12)
@@
Here, $\mathscr{C}(\mathcal{M}^{-})=\mathscr{C}(\mathcal{C})$, which is desirable when using (14) in set recursions involving the Pontryagin difference. However, solving (12) repeatedly can become computationally expensive for large ${N_{C}},{N_{S}}$. Also, it is unclear if such an approach extends to subtrahends beyond zonotopes, since it uses polytopic containment results from [23]. We address these shortcomings in this paper.
@@
where $\mathcal{M}^{-}$ is given in closed-form and $\mathscr{C}(\mathcal{M}^{-})=\mathscr{C}(\mathcal{C})$. Also, provide sufficient conditions for $\mathcal{M}^{-}=\mathcal{M}=\mathcal{M}^{+}$.
Our approach to address Prob. 1 follows a similar two-stage approach to that in [9] to obtain (14). However, unlike [9], we provide closed-form expressions for $\Gamma$ and $D$ used in (13) and (14) for a broader class of subtrahends that includes ellipsoids and convex unions of symmetric intervals. We also propose a constrained zonotopic outer-approximation to $\mathcal{M}$, and provide sufficient conditions under which these approximations are exact. The proposed outer-approximation relies on solving the following problem.
@@
We will also briefly discuss the relationship between the proposed approach and the two-stage approach in [9].